OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. csc 2x - 1

OpenStudy (anonymous):

Is it \[\csc ^{2}x-1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does this equal -cot

OpenStudy (anonymous):

1 + cot^2x = csc^2x right?

OpenStudy (anonymous):

is the answer cot^2x

OpenStudy (cggurumanjunath):



OpenStudy (cggurumanjunath):



OpenStudy (anonymous):

what's your doubt ?

OpenStudy (anonymous):

The ans is \[\cot ^{2}x\]

OpenStudy (anonymous):

following straight from one of the three basic trigonometric identities : \[\sin ^{2}x+\cos ^{2}x = 1\] \[\sec ^{2}x-\tan^{2}x = 1\] \[\csc^{2}x-\cot^{2}x = 1\]